A particle of mass M in circular motion on rough surface at an instant has velocity `v` and increases at rate `a`. What is the minimum value of "MU" or "U" . Got it for HC Verma. Answer is `a/g`

Grade:12

1 Answers

shwetank
9 years ago
Here the circular motion on rough surface is possible because of the force of friction, therefore
(U)Mg = Mv^2/r    (r = radius of circle)
(U) = v^2/rg
But v^2/r = a
So,
(U) = a/g